OpenStudy (anonymous):

Assume that y varies inversely as x. If x = 12 when y = 9, what is x when y = -3? possible answers- 108, -36, 27, -4

OpenStudy (anonymous):

y=kx 9=k(12) k=0.75 when y=-3 -3=kx -3/0.75 =x x=-4

OpenStudy (anonymous):

thank you :)